please help me to solve this question, the length of a rectangle is 2inches more than twice it's width. If the perimeter of the rectangle is 34 inches, find the dimensions of the retangle. I agains thank you in advance for your timely help
Answer by checkley71(8403)

PERIMETRER=2*LENGTH+2*WIDTH
LENGTH=2+2W THEREFORE THE PERIMETER IS
34=2(2+2W)+2W
34=4+4W+2W
34-4=6W
30=6W
W=30/6
W=5 THEN
LENGTH=2+2*5
LENGTH=2+10
LENGTH=12
PROOF
34=2*12=2*5
34=24+10
34=34